# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 4.1. Penrose tilings
Two important types of Penrose tilings arrowed rhombic Penrose tilings and the unmarked rhombic Penrose tilings (see [28], [29], [1]). The arrowed rhombic Penrose tiles, denoted $p_{5{\rm a}}$, are versions of the tiles from $p_{5}$ marked with arrows. Two of them are shown in Figure 2, (a). The entire set $p_{5{\rm a}}$ of $20$ prototiles is obtained by taking all rotations of these two tiles by multiples of $102\pi/$ 10. To make an FLC prototile set, the local rule $p^{(2)}_{5{\rm a}}$ requires two tiles to meet edge to edge with matching arrows (see Figure 2).
Figure 2. (a) two arrowed rhombic Penrose tiles, (b) arrowed rhombic Penrose tiling patch, and (c) the same patch with arrows erased. The erasure map $E$ has a local inverse $E^{-1}$.
Remark 4.1.
Penrose showed the following about $X_{p_{\text{5{\rm a}}}}$, (see e.g., [25] or [11]), although not expressing it in dynamical terminology: $X_{p_{\text{5{\rm a}}}}\neq\emptyset$ and every $x\in X_{p_{\text{5{\rm a}}}}$ is aperiodic, so $T$ is free. Moreover, $T$ minimal and uniquely ergodic (see [29]). These all properties essentially follow from the fact that there is a MLD topological conjugacy, called tiling substitution or inflation, between $X_{p_{\text{5{\rm a}}}}$ and the space $\gamma X_{p_{\text{5{\rm a}}}}$ of all tilings $x\in X_{p_{\text{5{\rm a}}}}$ expanded by $\gamma:=(1+\sqrt{5})/2$, (see e.g., [11], [29]).
The map $E:X_{p_{\text{5{\rm a}}}}\to X_{p_{5}}$ that erases the arrows is not onto (see [6]) since no periodic tiling can be corretly decorated with arrows. The set of unmarked Penrose tilings, defined by $\mathsf{X}:=E(X_{p_{\text{5{\rm a}}}})\subseteq X_{p_{5}}$ is a tiling space (see Figure 2 (c)), and we call the $(\mathsf{X},\mathbb{R}^{2},T)$ the (unmarked rhombic) Penrose tiling dynamical system.
@@ 4.2. The Kronecker model for Penrose tilings
Let $\mathbb{T}^{5}_{0}=\{{\boldsymbol{u}}\in\mathbb{T}^{5}:\{u_{0}+\cdots+u_{4}\}=0\}$, a closed subtorus of $\mathbb{T}^{5}$, homeomorphic to $\mathbb{T}^{4}$. For ${\boldsymbol{v}}_{j}=({\rm Re}(e^{2\pi i/5}),{\rm Im}(e^{2\pi i/5}))$, the $j$th 5th root of unity, let $W=[{\boldsymbol{v}}_{0}^{t};{\boldsymbol{v}}_{1}^{t};\dots;{\boldsymbol{v}}_{4}^{t}]$ be the $5\times 2$ matrix rows ${\boldsymbol{v}}_{j}$. Let $V$ be the $4\times 2$ matrix obtained by deleting the row of ${\boldsymbol{v}}_{3}^{t}$ from $W$. For ${\boldsymbol{t}}\in\mathbb{R}^{2}$, define $K^{\boldsymbol{t}}{\boldsymbol{u}}=\{{\boldsymbol{u}}+W{\boldsymbol{t}}\}$ on $\mathbb{T}^{5}$ and $L^{{\boldsymbol{t}}}{{\boldsymbol{u}}}=\{{{\boldsymbol{u}}}+V{{\boldsymbol{t}}}\}$ on $\mathbb{T}^{4}$. Clearly $(\mathbb{T}^{5},\mathbb{R}^{2},K)$ and $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ are Kronecker systems, with $\Sigma_{K}=\Sigma_{L}=V^{t}\mathbb{Z}^{5}=W^{t}\mathbb{Z}^{4}=\mathbb{Z}[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, the decomplexification of the complex ring generated by $e^{2\pi i/5}$. Note that $K$ is not minimal since $\mathbb{T}^{5}_{0}$ is $K$ invariant, but $L$ is minimal since $W^{t}$ is injective on $\mathbb{Z}^{4}$. Moreover, $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ is topologically conjugate to $(\mathbb{T}^{5}_{0},\mathbb{R}^{2},K|_{\mathbb{T}^{5}_{0}})$. This is the unique minimal Kronecker $\mathbb{R}^{2}$ dynamical system with $\Sigma_{K}=\mathbb{Z}[e^{2\pi i/5}]$.
The following is the main result of [29].
@@ 4.4. Duality and nonsingular tilings
For an edge to edge tiling $y$ of $\mathbb{R}^{2}$ by polygonal tiles, let $v(y)$ be its vertices, and $e(y)$ be its edges, with the tiles in $y$ being its faces. Another tiling $y^{*}$ by polygonal tiles is said to be geometrically dual to $y$ if there are bijections $y\overset{*}{\leftrightarrow}v(x)$, $v(y)\overset{*}{\leftrightarrow}x$ and $e(y)\overset{*}{\leftrightarrow}e(x)$, such that $\ell^{*}\in e(x)$, $\ell\in e(y)$ implies $\ell^{*}\perp\ell$.
We will now describe a geometric duality (see [6], [29]) between the grid tilings $y=y({\boldsymbol{u}})\in Y$ and a set of tilings $Y^{*}\supseteq X_{p5}$, described below. First, for a tile $E\in y({\boldsymbol{u}})$, define the vertex
@@
for arbitrary ${\boldsymbol{s}}\in{\rm int}(E)$. We connect two vertices $E_{1}^{*},E_{2}^{*}\in v(y^{*})$ by an edge $\ell^{*}\in e(y^{*})$ if their distance is $2/5$, or equivalently, if $E_{1},E_{2}\in y$ are adjacent across an edge $\ell\in e(y)$. The edges $e(y^{*})$ divide $\mathbb{R}^{2}$ into the polygonal tiles $D$ that make up $y^{*}$. In particular, if ${\boldsymbol{b}}\in v(y)$ is an $n$-fold crossing, then it is the common vertex of $2n$ tiles $E_{0},E_{1},\dots,E_{2n}\in y$, going clockwise. The duals $E^{*}_{0},E^{*}_{1},\dots,E^{*}_{2n}\in v(y^{*})$, and the connecting edges, enclose a $2n$-gon ${\boldsymbol{b}}^{*}:=D\in y^{*}$. The case $n=2$ is shown in Figure 5.
@@
We call a grid tiling $y=y({\boldsymbol{u}})\in Y$ nonsingular if every vertex ${\boldsymbol{b}}\in v(y)$ is a 2-fold crossing, and we denote the set of all nonsingular grid tilings by $Y_{1}$. Note that $y^{-1}(Y_{1})\subseteq\mathbb{T}^{5}_{0}$, or equivalently $Y_{1}\subseteq Y$, is dense $G_{\delta}$. Now suppose $y=y({\boldsymbol{u}})\in Y_{1}$, and the vertex ${\boldsymbol{b}}\in v(y)$ is a ($2$-fold) crossing of $\ell_{i,k}$ and $\ell_{j,k^{\prime}}$, $i<j$. Then its dual ${\boldsymbol{b}}^{*}$ in $y^{*}$ is a type ${\boldsymbol{v}}^{\prime}_{i}\wedge{\boldsymbol{v}}^{\prime}_{j}$ tile in $p_{5}$ (see Figure 5). Thus $y\in Y_{1}$ implies $y^{*}\in X_{p_{5}}$. But in fact, much more is true!
@@
Remark 4.7.
It is shown in [29] that $\varphi$ is uniformly continuous on $\mathsf{X}_{1}$, and $\mathsf{X}=\overline{\mathsf{X}_{1}}=\overline{Y^{*}_{1}}$. This is essentially deBruijn’s observation [6] that all Penrose tilings are limits of nonsingular Penrose tilings. The extension $\varphi:\mathsf{X}\to Y$ is the factor map in Theorem 4.3. The fact that $\varphi$ has a single preimage on the $G_{\delta}$ set $Y_{1}$ shows that $\varphi$ is almost 1:1.
@@ 4.5. The singular cases
If $y({\boldsymbol{u}})\notin Y_{1}$ it means that some of crossings in $y({\boldsymbol{u}})$ are not $2$-fold, and we call $y({\boldsymbol{u}})$ a singular grid tiling. Let $n$ be the highest crossing order in $y({\boldsymbol{u}})$. It is easy to see that ${\boldsymbol{u}}\in\mathbb{T}^{5}_{0}$ implies $n=4$ never occurs (see [6]). Thus there are two cases: Case A, $n=3$, which we will see corresponds to $Y_{2}$ in Theorem 4.3, and Case B, $n=5$, which corresponds to $Y_{10}$.
@@
The duals of $3$-fold crossings are hexagons, respectively wide or narrow. The hexagons line up in $y({\boldsymbol{u}})^{*}$ in configuration called an unfilled worm, Figure 6, (b), and we call the line $l$ along the center of the worm its axis. Since hexagons are not in $p_{5}$, the dual $y({\boldsymbol{u}})^{*}$ is not in $X_{p_{5}}$. But we can obtain a tiling in $X_{p_{5}}$ by “filling” each hexagon in $y({\boldsymbol{u}})^{*}$ by tiling it with three tiles from $p_{5}$. This can be done in one of two directions: direction $+$ will be when a single tile sits on the side facing ${\boldsymbol{v}}_{j}$, and direction -- will be when two tiles sit on the side facing ${\boldsymbol{v}}_{j}$
To obtain a Penrose tiling, all the hexagons must be filled in the same direction [6]. We denote the two resulting Penrose tilings as $x^{+}({\boldsymbol{u}}),x^{-}({\boldsymbol{u}})\in\mathsf{X}$, Figure 6, (c). The column of filled hexagons in $x^{+}({\boldsymbol{u}})$ or $x^{-}({\boldsymbol{u}})$ is called a filled worm. Going back and forth between $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ is called flipping the worm. Except for the tiles in the filled worm, all the tiles in $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ are the same. The following is the first part of the proof of Theorem 1.1.
@@
Suppose ${\mathcal{V}}={\rm span}\{{\boldsymbol{v}}_{j}^{\perp}\}$. Let $x^{+}\in\mathsf{X}_{2}$ and fix $r>0$. By applying a rotation and a translation we can assume that the single worm in $x^{+}$ is in direction ${\boldsymbol{v}}_{j}^{\perp}$ that lies outside ${\mathcal{V}}^{r}$. Let $x^{-}$ be $x^{+}$ with the worm flipped. Then $x^{+}[{\mathcal{V}}^{r}]=x^{-}[{\mathcal{V}}^{r}]$ but $x^{+}\neq x^{-}$. It follows from Definition 3.11 and Proposition 3.12 that $(\mathsf{X},\mathbb{R}^{2},T)$ is not expansive in the direction ${\mathcal{V}}$. ∎
@@
Case B: Here, $y=y({\boldsymbol{u}})$ has a single $5$-fold crossing at ${\boldsymbol{b}}\in\mathbb{R}^{2}$. By translation, we can assume ${\boldsymbol{b}}={\boldsymbol{0}}$ so the spines are $\ell_{j,0}$, $j=0,\dots,4$. All the crossings along these spines, except at ${\boldsymbol{0}}$, are wide and narrow $3$-fold, and all the other crossings in $y({\boldsymbol{u}})$ are $2$-fold (see [6]). In $y({\boldsymbol{u}})^{*}$ the dual of the $5$-fold crossing is a regular decagon, and there are unfilled half-worms of hexagons radiating from its faces. This configuration is called an unfilled cartwheel, and we call the direction ${\boldsymbol{v}}_{j}^{\perp}$ lines $l_{j}$ through the middle of each pair of half-worms spokes.

### v2

@@ 4.1. Penrose tilings
In the next two sections (4 and 5), we discuss two types of Penrose tilings: the arrowed rhombic Penrose tilings and the unmarked rhombic Penrose tilings (see [28], [29], [1]). Later, in Section 6, we will discuss two additional types of tilings whose tiling dynamical systems are topologically conjugate to the Penrose tiling dynamical system, (see Theorem 6.11).
The arrowed rhombic Penrose tiles, denoted by $p_{5{\rm a}}$, are versions of the tiles from $p_{5}$ (Example 3.6) marked with arrows. Two of these tiles are shown in Figure 2(a). The entire set $p_{5{\rm a}}$ of $20$ prototiles is obtained by taking all rotations of these two tiles by multiples of $102\pi/$ 10. To make an FLC prototile set, the local rule $p^{(2)}_{5{\rm a}}$ requires two tiles to meet edge to edge with matching arrows (see Figure 2).
Figure 2. (a) two arrowed rhombic Penrose tiles, (b) arrowed rhombic Penrose tiling patch, and (c) the same patch with arrows erased. The erasing map $E$ has a local inverse $E^{-1}$.
The erasing map $E:X_{p_{\text{5{\rm a}}}}\to X_{p_{5}}$, which erases the arrows, is not onto (see [6]) since no periodic tiling can be correctly decorated with arrows. Clearly $E(X_{p_{\text{5{\rm a}}}})$ is a tiling space.
Definition 4.1.
The unmarked Penrose tiling tiling space is defined by $\mathsf{X}:=E(X_{p_{\text{5{\rm a}}}})\subseteq X_{p_{5}}$, and we call the $(\mathsf{X},\mathbb{R}^{2},T)$ the (unmarked rhombic) Penrose tiling dynamical system.
@@ 4.2. The Kronecker model for Penrose tilings
Let $\mathbb{T}^{5}_{0}=\{{\boldsymbol{u}}\in\mathbb{T}^{5}:\{u_{0}+\cdots+u_{4}\}=0\}$, a closed subtorus of $\mathbb{T}^{5}$, homeomorphic to $\mathbb{T}^{4}$. For ${\boldsymbol{v}}_{j}=({\rm Re}(e^{2\pi i/5}),{\rm Im}(e^{2\pi i/5}))$, the $j$th 5th root of unity, let $W=[{\boldsymbol{v}}_{0}^{t};{\boldsymbol{v}}_{1}^{t};\dots;{\boldsymbol{v}}_{4}^{t}]$ be the $5\times 2$ matrix rows ${\boldsymbol{v}}_{j}$. Let $V$ be the $4\times 2$ matrix obtained by deleting the row of ${\boldsymbol{v}}_{3}^{t}$ from $W$. For ${\boldsymbol{t}}\in\mathbb{R}^{2}$, define $K^{\boldsymbol{t}}{\boldsymbol{u}}=\{{\boldsymbol{u}}+W{\boldsymbol{t}}\}$ on $\mathbb{T}^{5}$ and $L^{{\boldsymbol{t}}}{{\boldsymbol{u}}}=\{{{\boldsymbol{u}}}+V{{\boldsymbol{t}}}\}$ on $\mathbb{T}^{4}$. Clearly $(\mathbb{T}^{5},\mathbb{R}^{2},K)$ and $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ are Kronecker systems, with $\Sigma_{K}=\Sigma_{L}=V^{t}\mathbb{Z}^{5}=W^{t}\mathbb{Z}^{4}=\mathbb{Z}[e^{2\pi i/5}]\subseteq\mathbb{R}^{2}$, the decomplexification of the complex ring generated by $e^{2\pi i/5}$. Note that $K$ is not minimal since $\mathbb{T}^{5}_{0}$ is $K$ invariant, but $L$ is minimal since $W^{t}$ is injective on $\mathbb{Z}^{4}$. Moreover, $(\mathbb{T}^{4},\mathbb{R}^{2},L)$ is topologically conjugate to $(\mathbb{T}^{5}_{0},\mathbb{R}^{2},K|_{\mathbb{T}^{5}_{0}})$. This is the unique minimal Kronecker $\mathbb{R}^{2}$ dynamical system with $\Sigma_{K}=\mathbb{Z}[e^{2\pi i/5}]$. The following is the main result of [29].
@@
Previously, Penrose [25] had obtained the basic properties of Penrose tilings, (many of which also follow from Theorem 4.3), although not expressing his results in dynamical terminology. He showed that every $x\in X_{p_{\text{5{\rm a}}}}$ is aperiodic, so $T$ is free, and also that $T$ is minimal and uniquely ergodic (see [29]). Penrose’s proofs of these property are based on the following remarkable scaling property for Penrose tilings.
Proposition 4.4 (Penrose, [25], see also [11], [28]).
There is a MLD topological conjugacy (see Definition 3.5 and the comments that follow it) $C:X_{p_{\text{5}}}\to X_{(1/\gamma)p_{\text{5}}}$ (or $C:\mathsf{X}\to(1/\gamma)\mathsf{X}$ ), where $(1/\gamma)p_{\text{5{\rm a}}}$ (or $(1/\gamma)\mathsf{X}$) is same set of tilings, scaled down by $1/\gamma$, (see [29] for a picture of $C$).
In addition, the map $C^{\prime}:=\gamma C:X_{p_{\text{5{\rm a}}}}\to X_{p_{\text{5{\rm a}}}}$, or $C^{\prime}:\mathsf{X}\to\mathsf{X}$, (which scales back up to the original scale) called a tiling substitution or inflation, is also a topological conjugacy (see [28]). Note that in [1] this would be called an improper inflation as opposed to a stone inflation, because the map $C$ does not perfectly tile each $D\in p_{\text{5}}$ with tiles $D^{\prime}\in(1/\gamma)p_{\text{5{\rm a}}}$. The scaled map $c^{\prime}=\gamma c$ corresponding to $C^{\prime}:=\gamma C$ (see Remark 3.5) can be applied iteratively to tiling patches rather than tiles, allowing one to construct ever larger patches thereby proving $X_{p_{\text{5{\rm a}}}}\neq\emptyset$.
@@ 4.4. Duality and nonsingular tilings
For an edge to edge tiling $y$ of $\mathbb{R}^{2}$ by polygonal tiles, let $v(y)$ be its vertices, and $e(y)$ be its edges, with the tiles in $y$ being its faces. Another edge to edge tiling $y^{*}$ by polygonal tiles is said to be geometrically dual to $y$ if there are bijections $y\overset{*}{\leftrightarrow}v(x)$, $v(y)\overset{*}{\leftrightarrow}x$ and $e(y)\overset{*}{\leftrightarrow}e(x)$, such that $\ell^{*}\in e(x)$, $\ell\in e(y)$ implies $\ell^{*}\perp\ell$.
We will now describe a geometric duality (see [6], [29] for more details) between the grid tilings $y=y({\boldsymbol{u}})\in Y$ and a certain set of tilings $Y^{*}\supseteq X_{p5}$. First, for a tile $E\in y({\boldsymbol{u}})$, we define the vertex
@@
for arbitrary ${\boldsymbol{s}}\in{\rm int}(E)$. We connect two vertices $E_{1}^{*},E_{2}^{*}\in v(y^{*})$ by an edge $\ell^{*}\in e(y^{*})$ if their distance is $2/5$, or equivalently, if $E_{1},E_{2}\in y$ are adjacent across an edge $\ell\in e(y)$. The edges $e(y^{*})$ divide $\mathbb{R}^{2}$ into the polygonal tiles $D$ that make up $y^{*}$. In particular, if ${\boldsymbol{b}}\in v(y)$ is an $n$-fold crossing, then it is the common vertex of $2n$ tiles $E_{0},E_{1},\dots,E_{2n}\in y$, going clockwise. The duals $E^{*}_{0},E^{*}_{1},\dots,E^{*}_{2n}\in v(y^{*})$ and the edges in $e(y^{*})$ connecting them enclose a $2n$-gon ${\boldsymbol{b}}^{*}:=D\in y^{*}$. The case $n=2$ is shown in Figure 5.
@@
We call a grid tiling $y=y({\boldsymbol{u}})\in Y$ nonsingular if every vertex ${\boldsymbol{b}}\in v(y)$ is a 2-fold crossing. We denote the set of all nonsingular grid tilings by $Y_{1}$. Note that $y^{-1}(Y_{1})\subseteq\mathbb{T}^{5}_{0}$ is dense $G_{\delta}$, as is $Y_{1}\subseteq Y$. Now suppose $y=y({\boldsymbol{u}})\in Y_{1}$, and the vertex ${\boldsymbol{b}}\in v(y)$ is a ($2$-fold) crossing of $\ell_{i,k}$ and $\ell_{j,k^{\prime}}$, $i<j$. Then its dual ${\boldsymbol{b}}^{*}$ in $y^{*}$ is a type ${\boldsymbol{v}}^{\prime}_{i}\wedge{\boldsymbol{v}}^{\prime}_{j}$ tile in $p_{5}$ (see Figure 5). Thus $y\in Y_{1}$ implies $y^{*}\in X_{p_{5}}$. But in fact, much more is true.
@@
Lemma 4.8 (See [29]).
The map $\varphi:\mathsf{X}_{1}\to Y_{1}$ is uniformly continuous and has a continuous extension to $\varphi:\mathsf{X}\to Y$ where $\mathsf{X}:=\overline{\mathsf{X}_{1}}$.
This is essentially deBruijn’s observation [6] that all Penrose tilings are limits of nonsingular Penrose tilings. The extension $\varphi:\mathsf{X}\to Y$ is the factor map in Theorem 4.3. The fact that $\varphi$ has a single preimage on the $G_{\delta}$ set $Y_{1}$ shows that $\varphi$ is almost 1:1.
@@ 4.5. The singular cases
If $y({\boldsymbol{u}})\notin Y_{1}$ it means that some of crossings in $y({\boldsymbol{u}})$ are not $2$-fold. We call $y({\boldsymbol{u}})$ a singular grid tiling. Let $n$ be the highest crossing order in $y({\boldsymbol{u}})$. It is easy to see that ${\boldsymbol{u}}\in\mathbb{T}^{5}_{0}$ implies that $n=4$ never occurs (see [6]). Thus there are two cases: Case A, $n=3$, which we will see corresponds to $Y_{2}$ in Theorem 4.3, and Case B, $n=5$, which corresponds to $Y_{10}$.
@@
The duals of $3$-fold crossings are hexagons, respectively wide or narrow. The hexagons line up in $y({\boldsymbol{u}})^{*}$ in configuration called an unfilled worm, Figure 6(b), and we call the line $l$ along the center of the worm its axis. Since hexagons are not in $p_{5}$, the dual $y({\boldsymbol{u}})^{*}$ is not in $X_{p_{5}}$. We can obtain a tiling in $X_{p_{5}}$ by “filling” each hexagon in $y({\boldsymbol{u}})^{*}$ by tiling it with three tiles from $p_{5}$. This can be done in one of two directions: direction $+$, meaning a single tile sits on the side facing ${\boldsymbol{v}}_{j}$, and direction --, meaning two tiles sit on the side facing ${\boldsymbol{v}}_{j}$, (see Figure 6(b) and (c)).
To obtain a Penrose tiling, all the hexagons must be filled in the same direction [6]. We denote the two resulting Penrose tilings as $x^{+}({\boldsymbol{u}}),x^{-}({\boldsymbol{u}})\in\mathsf{X}$, Figure 6(b) and (c). The column of filled hexagons in $x^{+}({\boldsymbol{u}})$ or $x^{-}({\boldsymbol{u}})$ is called a filled worm. Going back and forth between $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ is called flipping the worm. Except for the tiles in the filled worm, all the tiles in $x^{+}({\boldsymbol{u}})$ and $x^{-}({\boldsymbol{u}})$ are the same. The following is the first part of the proof of Theorem 1.1.
@@
Suppose ${\mathcal{V}}=\{{\boldsymbol{v}}_{j}\}^{\perp}$. Let $x^{+}\in\mathsf{X}_{2}$ and fix $r>0$. By applying a rotation and a translation we can assume that the single worm in $x^{+}$ is in a direction ${\boldsymbol{v}}_{j}^{\perp}$ that lies outside ${\mathcal{V}}^{r}$. Let $x^{-}$ be $x^{+}$ with the worm flipped. Then $x^{+}[{\mathcal{V}}^{r}]=x^{-}[{\mathcal{V}}^{r}]$ but $x^{+}\neq x^{-}$. It follows from Definition 3.12 and Proposition 3.13 that $(\mathsf{X},\mathbb{R}^{2},T)$ is not expansive in the direction ${\mathcal{V}}$. ∎
@@
Case B: Here, $y=y({\boldsymbol{u}})$ has a single $5$-fold crossing at ${\boldsymbol{b}}\in\mathbb{R}^{2}$. By translation, we can assume ${\boldsymbol{b}}={\boldsymbol{0}}$ so the spines are $\ell_{j,0}$, $j=0,\dots,4$. All the crossings along these spines, except at ${\boldsymbol{0}}$, are wide and narrow $3$-fold, and all the other crossings in $y({\boldsymbol{u}})$ are $2$-fold (see [6]). In $y({\boldsymbol{u}})^{*}$ the dual of the $5$-fold crossing is a regular decagon, and there are unfilled half-worms of hexagons radiating from its faces. This configuration is called an unfilled cartwheel, and we call a direction ${\boldsymbol{v}}_{j}^{\perp}$ line $l_{j}$ a spoke if it goes through the middle of a pair of half-worms.